OpenStudy (anonymous):

The drawing of a wheel in a book is done at one-sixteenth scale. If the drawing is 1.8 inches in diameter, what is the wheel's diameter? A. 32" B. 28.8" C. 24" D. 0.1125" @nincompoop

OpenStudy (anonymous):

If 1.8 is 1/16 of the actual diameter, you would want to multiply 1.8 by 16 to get the full scale measurement

OpenStudy (anonymous):

you would just augment or multiply by 16 your model's 1.8 diameter ...

OpenStudy (anonymous):

I don't get this because of 1/16. Why do have to multiply?

OpenStudy (anonymous):

If 1.8 is 1/16 of a whole, it means it is 1 part out of 16, meaning you would originally divide the full scale by 16 parts to determine what 1 of the 16 parts is, knowing this you want to work backwards since you already have the 1/16 determined, and multiply it by 16 to in turn receive the full amount of the original scale

OpenStudy (nincompoop):

@zJicez is correct 1/16x = 1.8 16 × 1/16x = 1.8 × 16 x = 1.8 × 16

OpenStudy (nincompoop):

the "ex" and multiplication symbol look almost the same >.< the latter is wider though if you look at it closely.

OpenStudy (nincompoop):

does the algebra break down explain it intuitively why you should multiply by 16 @legendarymath ?

OpenStudy (nincompoop):

similarly when you have 4x = 20 you divide both sides by 4 to obtain x alone 4x ÷ 4 = 20 ÷ 4 of course whatever you do on one side of the equation, must be done on the other to keep the equation balanced. x = 5

OpenStudy (nincompoop):